# and quantum information science

### v1

@@ 7.5.1 Spread complexity in double-scaled SYK model equals wormhole length
To compute the spread complexity of the DSSYK model we need the Lanczos coefficients. As we discussed above these can be computed either by tridiagonalizing the Hamiltonian by the moment method, or through an integral formula applied to the density of states. To apply the moment method we have to calculate the expectation values of powers of the Hamiltonian in the initial state which we take to be the thermo-field double. So we have to calculate
$M_{2k}=\langle\mathrm{Tr}(H^{2k})\rangle=i^{kp}\sum_{I_{1}\cdots I_{2k}}\langle J_{I_{i}}\cdots J_{2k}\rangle\mathrm{Tr}(\psi_{I_{1}}\cdots\psi_{I_{n}})\,,$ (119)
@@
where we introduced the multi-index $I_{k}=(i_{ik}\cdots i_{pk})$. We have only written the even moments because the odd moments vanish. To finish the computation we have to Wick contract the multi-indexed fermion terms in pairs. As discussed in [317, 318] the Feynman diagrams describing these contractions are called chord diagrams and can be visualized by placing 2k points on the circumference of a circle and then contracting the dots in pairs. Carrying out these contractions we arrive at the moments
@@
$M_{2k}=(1-q)^{k}\mathcal{J}^{2k}\sum_{\text{diag. with k chords}}q^{\text{number of intersections}}=\langle 0|T^{2k}|0\rangle\,.$ (120)
@@
The second equality expresses the moments in terms of a triagonal transfer matrix $T$ acting on $\lvert 0\rangle$, an auxiliary state with zero chords, that represents the thermofield double state at infinite temperature [318]. Essentially by construction $T$ is the tridiagonalized Hamiltonian, the apparently auxiliary chord basis is precisely the Krylov basis. Now, following the moment method, one can automatically read-off Lanczos coefficients. They are
$b_{n}=J\sqrt{\frac{1-q^{n}}{\lambda(1-q)}}\,.$ (121)
@@
We can then solve the Schrodinger equation, and compute spread complexity. In the double scaling limit, $N\to\infty$ and hence the Hilbert space is infinite dimensional. Using the Lanczos coefficients in (121) we find a spread complexity that grows linearly forever. Meanwhile we can apply an additional low-energy scaling limit [318]. This so-called triple scaling limit is effectively a continuum limit for the chord basis in which the $T$, the tridiagonalized Hamiltonian, reduces to a boundary description of a bulk Jackiw-Teitelboim (JT) gravity. In this limit, [316, 26] showed that the quantity calculating the length of the wormhole in the TFD state of JT gravity is precisely the spread complexity being computed above. This construction establishes a precise correspondence between wormhole length in gravity and spread complexity in this limit of JT gravity and the dual DSSYK model.
@@
However, this construction leads to a puzzle. At any finite $N$, the Hilbert space should be finite dimensional and hence the spread complexity should saturate at late times, and along with it the quantum analog of wormhole length. How can we see this? Fortunately, the results of [317] demonstrate that these moments of the SYK Hamiltonian described above arise from the distribution
@@
In other words this is the density of states at the given $q$. Following the discussion in Sec. 6.1.4 we can directly compute the Lanczos spectrum from this density of states via an integral formula. This procedure was carried out in [27], where the authors demonstrated that the the Lanczos coefficients for the SYK model match the computations of [318] described above in an initial range of Krylov indices, but then decline systematically to zero. The decline occurs over the whole range of the Hilbert space, with a slope of $1/N$, where $N$ is the dimension of the Hilbert space. This descent to zero for the Lanczos coefficients causes the spread complexity to saturate at late times. If we maintain the identification between wormhole length and spread complexity that applies at early times, we would say that the wormhole length is saturating at late times. From the bulk JT gravity point of view this decline thus corresponds to a non-perturbative quantum effect in gravity that corrects the steady increase in the Lanczos coefficients and wormhole size expected from the early time classical description of the system. Of course at late times we could imagine alternate non-perturbative definitions of the wormhole length in JT gravity that also agree with the early time identification with spread complexity. Interesting developments along this line are presented in [319, 225].
@@
In a related vein, [320] utilized new developments identifying the gravity dual to DSSYK [321] to show that the relation between the spread complexity and the bulk volume extends past the limits used in the studies of [26] to the full DSSYK regime at arbitrary temperatures. The key finding of [320] was that the match with the volume on the gravity side requires computing spread complexity for the finite temperature state preparation using the Euclidean path integral, as opposed to including the Euclidean state preparation in the reference state and assigning complexity only to purely Lorentzian evolution. Furthermore, as the calculations in [320] extended also to the quantum regime, they were able to extract the form of the leading correction to the volume=complexity proposal originating from the contributions of quantum fields in two dimensional bulk. This correction was crucially needed for the quantitative match with the boundary spread complexity. Interestingly, the correction exhibits behavior compatible with it being the spread complexity of the bulk quantum fields, but a precise relation remains unknown.
@@
These findings suggest that Krylov-basis definitions of complexity are natural from a bulk perspective and could provide a non-perturbative definition of bulk volumes. A crucial next step would be to extend these studies to higher dimensions, beginning with AdS3. While the powerful exact results available in DSSYK may not extend directly, recent progress with Virasoro TQFTs [322] could offer insights into the bulk Hilbert space and its relation to the Krylov basis.
@@ 7.5.2 Rate of spread-complexity growth and proper momentum
The above comparisons between the Krylov basis definitions of complexity and holography were made in the context of low-dimensional toy models. This progress was possible due to the explicit definition of the thermofield-double state in the double-scaled SYK model using the transfer matrix (and its $q$–deformed algebra). However, extending this approach to higher-dimensional, genuinely holographic setups may be significantly more challenging. Nevertheless, several universal insights about operator growth and spread complexity can already be tested in higher dimensions. One of the most concrete examples is the relation between the growth rate of operator complexity (or size) and the radial momentum of its dual massive particle in AdS. Intuitively, we expect that the time derivative of complexity (the rate of complexity) of the Heisenberg operator $\mathcal{O}(t)$ should be proportional to the radial momentum of the particle
@@
This relation has been argued in various ways in [38, 323, 324, 120, 325], including qualitative models of operator growth within epidemic models [326]. Furthermore, Refs. [327, 328, 329] essentially proved this relation under the assumption of the CV proposal. However, at that time, a precise definition of complexity satisfying Eq. (123) was missing.
By now, there are several setups where this formula can be precisely evaluated on both sides of the holographic correspondence. Firstly, in the SYK model at large $q$, one can directly compute the expectation value of the size operator [145, 330], which is defined in terms of two copies of fermions (left and right in the thermofield-double construction, denoted with subscripts $l$ and $r$ respectively) as
@@
and match its rate with the radial momentum of a particle in AdS2 [330] (see also [331]).
More recently, this correspondence was also confirmed between the spread complexity in 2D CFTs and the proper radial momentum of particles in AdS3 spacetime. Reference [332] studied a local operator quench where a CFT state is locally excited by a primary operator with conformal dimension $\Delta$, and evolved unitarily with the Hamiltonian. Holographically, this setting corresponds to a particle with mass $m=\Delta$ propagating from the boundary towards the bulk, with the particle’s initial position corresponding to the operator’s energy regulator $\epsilon$. Interestingly, the rate of spread complexity for this dynamical state is related to the particle’s momentum via
$\partial_{t}\mathcal{C}_{K}(t)=-\frac{1}{\epsilon}P_{\rho}(t).$ (125)
However, the key observation is that the momentum must be computed in the proper radial distance coordinate $\rho$ (for example, this relation would not hold in Poincaré coordinates). The relation between Krylov complexity and momentum has also been analyzed from a slightly different perspective in [333, 334]. The connection between the position on the Krylov chain $n$ and (proper) radial distance in the bulk was also substantiated in [335].
Finally, the evolution of spread complexity following local quench protocols in 2D CFTs was studied in [136]. These more geometric setups also follow the pattern of the SL(2,R) analytical solutions discussed in Sec. 6.3.1. They studied the case of joining quenches where, at $t=0$, two uncoupled CFTs (on half-spaces) at finite temperature were joined at $x=0$ and then evolve unitarily. In this context, the spread complexity was proportional to the central charge of the CFT as follows
@@
This local quench setup is described holographically using the AdS/BCFT correspondence [336], where the end-of-the-world (EOW) brane is time-dependent. Interestingly, the tip of the EOW brane, which probes the deepest part of the AdS bulk, follows a geodesic of a massive particle (with heavy mass $m=c/16$). The radial momentum of this particle, computed in the proper distance coordinate, satisfies
$\partial_{t}\mathcal{C}_{K}(t)=-\frac{1}{2\epsilon}P_{\rho}(t).$ (127)
These explicit checks suggest that, depending on the initial state, it may be possible to engineer precise holographic setups where the spread complexity rate can be directly compared to a quantity in gravity (and integrated). Particularly encouraging is the overall dependence on the central charge in (126), which was also crucial in comparing the RT formula [337] with entanglement entropy in 2D CFTs. To finish, let us point that an intriguing perspective on the Krylov basis was also presented in [137], where, using Wigner functions, the authors argued that the Krylov basis is “the most classical" basis in QFT, potentially making it a natural candidate for describing the evolution of semi-classical states in the bulk code subspace.
7.5.3 Nielsen’s complexity and geometry
There are also attempts to understand holographic complexity by exploiting equivalence between bulk and boundary symplectic forms [338, 36, 24], inspired by behavior under conformal transformations in Bañados geometries [118, 339].
The relation to symplectic forms also motivated the connection between quantum complexity in CFTs and gravity for the FS cost function obtained in section 5.5.3. To understand this relation, we observe that one can map trajectories built from the unitary operators (46) of a CFTd to the geodesics of a massive particle of mass $m$ in AdSd+1 spacetime. The FS metric (47a) receives a natural geometric interpretation in terms of the minimal $\delta X_{\rm min}$ and maximal $\delta X_{\rm max}$ perpendicular distances between infinitesimally nearby geodesics (as illustrated in Fig. 1 of reference [32])
@@ 7.5.3 Nielsen’s complexity and geometry
It would be very interesting to generalize the relation (128) to complexity associated with states that are related by a finite transformation. Furthermore, the connection to other than Fubini-Study cost functions is a subject of an ongoing investigation [121].
We should stress that the considerations in Sec. 5.5.3 and the relation (128) belong to the case (a) in Fig. 6. Also, the bulk quantity defined by Eq. (128) is special to this setup in which different states are parametrized by geodesic trajectories in the empty AdS. In this special situation it makes sense to put two such trajectories together in the same AdS geometry. For general excited states in holography one deals with two distinct bulk geometries whose direct comparison occurs via covariant notions (for example, through one of the holographic complexity proposals) associated with each of the corresponding asymptotic boundaries.
@@
As a result, a natural question in this context is about the generality of the relation (128). In the case (a) in Fig. 6 it is unknown. However, the case (b) in Fig. 6 is different and allows in fact to make a general statement, as originally proposed in [34]. The key idea is the following. For circuits in which the circuit parameter is the physical time in QFT, one layer of the circuit is simply given by the instantaneous QFT Hamiltonian. This Hamiltonian is defined in terms of the energy-momentum tensor components smeared along a time slice, e.g. $H=\int d^{d-1}x\,T_{tt}(\vec{x})$. The infinitesimal increment of the Fubini-Study distance (47a) is proportional to the variance of the Hamiltonian in an instantaneous state. However, given that the Hamiltonian is a sum of local operators, its variance is given by a sum of two-point functions of local operators. In AdS/CFT, the holographic dictionary [313, 314, 340, 341] provides a systematic geometric way of computing correlations functions of the boundary QFTs. As a result, the Fubini-Study cost (47a) associated with one layer of time evolution on the boundary is always geometric by the virtue that all its ingredients are geometric. This statement therefore expresses an exact mapping between two geometries: the auxiliary complexity geometry of the Fubini-Study cost on one hand and the gravitational geometry of the bulk.
It is important to stress that while the above statement is geometric, the originating bulk manifestation of the Fubini-Study cost is rather complicated and implicit. However, in the context of holographic two-dimensional CFT and states represented holographically as solutions of the Einstein’s equations with negative cosmological constant in three bulk dimensions one can relate two-point functions of the energy-momentum tensor to geodesic lengths in the bulk [34]. This leads to a much more explicit expression for the Fubini-Study cost and a direct connection to the kinematic space program [342, 343, 344].
At the moment of writing this review there are two major open holographic questions relevant for the Nielsen’s complexity. The first one has to do with a holographic interpretation of other than the Fubini-Study cost functions. This question is a question of infinitesimal transformations. For finite transformations, the major open question is about even a single explicit holographic manifestation of Nielsen’s complexity. The goal when considering these problems should be to find a way of thinking about holography and Nielsen’s complexity that would allow for a general proof of what the gravitational manifestation of the latter is. However, it is also possible that such a relation does not exist in a simple form. Indeed, the results described in previous subsections indicate that perhaps one should seek for the microscopic explanation of some members of holographic complexity proposals in terms of the Krylov (spread) complexity.

### v2

@@ 7.5.1 Spread complexity in double-scaled SYK model equals wormhole length
To compute the spread complexity of the DSSYK model we need the Lanczos coefficients. As we discussed, these can be computed either by tridiagonalizing the Hamiltonian by the moment method, or through an integral formula applied to the density of states. To apply the moment method we have to calculate the expectation values of powers of the Hamiltonian in the initial state which we take to be the thermo-field double state. So we have to calculate
$M_{2k}=\langle\mathrm{Tr}(H^{2k})\rangle=i^{kp}\sum_{I_{1}\cdots I_{2k}}\langle J_{I_{1}}\cdots J_{I_{2k}}\rangle\mathrm{Tr}(\psi_{I_{1}}\cdots\psi_{I_{2k}})\,,$ (130)
@@
where we introduced the notation $\psi_{I}=\psi_{i_{1}}\psi_{i_{2}}\cdots\psi_{i_{p}}$. We have only written the even moments because the odd moments vanish. To finish the computation we have to Wick contract the multi-indexed fermion terms in pairs. As discussed in [368, 32], the Feynman diagrams describing these contractions are called chord diagrams and can be visualized by placing $2k$ points on the circumference of a circle, and then contracting the dots in pairs. Carrying out these contractions, we arrive at the moments
@@
$M_{2k}=(1-q)^{k}\mathcal{J}^{2k}\sum_{\text{diag. with k chords}}q^{\text{number of intersections}}=\langle 0|T^{2k}|0\rangle\,,$ (131)
@@
where $q=e^{-\lambda}$ The second equality expresses the moments in terms of a tridiagonal transfer matrix $T$ acting on $\lvert 0\rangle$, an auxiliary state with zero chords, that represents the thermofield double state at infinite temperature [32]. Essentially, by construction $T$ is the tridiagonalized Hamiltonian; the apparently auxiliary chord basis is precisely the Krylov basis of Sec. 6. Now, following the moment method, we can immediately read off the Lanczos coefficients. They are
$b_{n}=\mathcal{J}\sqrt{1-q^{n}}\,.$ (132)
@@
We can then solve the Schrodinger equation, and compute spread complexity. In the double scaling limit $N\to\infty$, and hence the Hilbert space is infinite-dimensional. Correspondingly, using the Lanczos coefficients in (132), we find a spread complexity that grows linearly forever (see Sec. 6). Meanwhile, we can apply an additional low-energy scaling limit [32]. This so-called triple scaling limit is effectively a continuum limit for the chord basis in which $T$, the tridiagonalized Hamiltonian, reduces to a boundary description of a bulk Jackiw-Teitelboim (JT) gravity. In this limit, the authors of [33, 34] showed that the quantity calculating the length of the wormhole in the TFD state of JT gravity is precisely the spread complexity discussed above. This construction establishes a precise correspondence between wormhole length in gravity and spread complexity in this limit of JT gravity and the dual DSSYK model.
@@
However, this construction leads to a puzzle. At any finite $N$, the Hilbert space should be finite-dimensional, and hence the spread complexity should saturate at late times, together with the quantum analog of wormhole length. How can we see this? Fortunately, the results of [368] demonstrate that the moments of the SYK Hamiltonian described above arise from the distribution
@@
In other words, this is the density of states at a given $q$. Following the discussion in Sec. 6.1.4, we can directly compute the Lanczos spectrum from this density of states at large but finite $N$ via an integral formula. This procedure was carried out in [249, 35], where the authors demonstrated that the Lanczos coefficients for the SYK model match the computations of [32] (described above) for an initial range of Krylov indices, but then decline systematically to zero. The decline occurs over the whole range of Krylov indices up to the dimension of the Hilbert space. with a slope of $1/N$, where $N$ is the dimension of the Hilbert space. This descent to zero for the Lanczos coefficients causes the spread complexity to saturate at late times. If we maintain the identification between wormhole length and spread complexity that applies at early times, we would say that the wormhole length is saturating at late times. From the bulk JT gravity point of view, this decline thus corresponds to a non-perturbative quantum effect in gravity that corrects the steady increase in the Lanczos coefficients and wormhole size expected from the early-time classical description of the system. Of course, at late times we could imagine alternative non-perturbative definitions of the wormhole length in JT gravity that also agree with the early-time identification with spread complexity. Interesting recent developments along this line are also presented in [31, 369, 269] which were discussed in Sec. 7.2.
@@
In a related vein, Ref. [370] utilized new developments identifying the gravity dual of DSSYK [371] to show that the relation between spread complexity and the bulk volume extends past the limits used in the studies of [34]. In particular, this relation extends to the full DSSYK regime at arbitrary temperatures. The key finding of [370] was that the match with the volume on the gravity side required computing the spread complexity for the preparation of a finite-temperature state using the Euclidean path integral. This is opposed to including the Euclidean state preparation in the reference state, and then assigning non-vanishing complexity only to the purely Lorentzian evolution. Furthermore, since the calculations in [370] also applied to the quantum regime, the authors were able to extract the form of the leading correction to the CV proposal originating from the contributions of quantum fields in a two-dimensional bulk geometry. This correction was crucially needed for the quantitative match with the boundary spread complexity. Interestingly, the correction exhibits a behavior compatible with the expectations for the spread complexity of bulk quantum fields, but a precise relation remains unknown.
@@
These findings suggest that Krylov-basis definitions of complexity are natural from a bulk perspective, and could provide a non-perturbative definition of bulk volumes. A crucial next step would be to extend these studies to higher dimensions, beginning with AdS3. While the powerful exact results available in DSSYK may not extend directly, recent progress with Virasoro TQFTs [372] could offer insights into the bulk Hilbert space and its relation to the Krylov basis. Another approach could be to exploit the observation in [373] that at late times the volume of an extremal slice through the ERB of any spherically-symmetric black hole reduces to a geodesic length in an effective two-dimensional JT gravity theory. This suggests that we may be able to directly export the results relating spread complexity and wormholes in JT gravity to higher dimensions.
@@ 7.5.2 Rate of spread-complexity growth and proper momentum
The above comparisons between the Krylov basis definitions of complexity and holography were made in the context of low-dimensional toy models. This progress was possible due to the explicit definition of the thermofield-double state in the double-scaled SYK model using a transfer matrix and its $q$–deformed algebra. However, extending this approach to higher-dimensional holographic setups may be challenging. Nevertheless, several universal insights about operator growth and spread complexity can already be tested in higher dimensions. One of the most concrete examples is the relation between the growth rate of operator complexity (or size) and the radial momentum of its dual massive particle in AdS spacetime. Intuitively, we expect that the time derivative of complexity (the rate of complexity) of the Heisenberg operator $\mathcal{O}(t)$ should be proportional to the radial momentum of the particle
@@
Since $\mathcal{C}$ is the mean position in the Krylov chain, this relation is telling us that the rate of progress along this chain is dual to the radial momentum. This sort of relation has been argued for in various ways in [49, 374, 375, 376, 248], including qualitative models of operator growth using epidemic models [377]. Furthermore, Refs. [378, 379, 380] essentially proved this relation under the assumption of the CV proposal. However, at that time, a precise definition of complexity satisfying Eq. (134) was missing.
By now, there are setups where this formula can be precisely evaluated on both sides of the holographic correspondence. Firstly, in the SYK model at large $q$, one can directly compute the expectation value of the size operator [165, 381] that can be defined in terms of two copies of fermions (left and right in the thermofield-double construction, denoted with subscripts $l$ and $r$ respectively) as
@@
and in [381], authors showed that its rate of change in time (i.e., time derivative) is proportional to the radial momentum of a particle in AdS2 (see also [382]).
Recently, this correspondence was also confirmed between the spread complexity in 2D CFTs and the proper radial momentum of particles in AdS3 spacetime. Reference [383] studied a local operator quench where a CFT state is locally excited by a primary operator with conformal dimension $\Delta$, and evolved unitarily with the Hamiltonian. Holographically, this setting corresponds to a particle with mass $m=\Delta$ propagating from the boundary towards the bulk, with the particle’s initial position corresponding to the operator’s energy regulator $\varepsilon$. Interestingly, the rate of spread complexity for this dynamical state is related to the particle’s momentum via
$\partial_{t}\mathcal{C}_{K}(t)=-\frac{1}{\varepsilon}P_{\rho}(t).$ (136)
However, the key observation is that the momentum must be computed in the proper radial distance coordinate $\rho$ (for example, this relation would not hold in Poincaré coordinates). This relation between Krylov complexity and momentum has also been analyzed from a slightly different perspective in [384, 385], and the connection between the position on the Krylov chain $n$ and proper radial distance in the bulk was further substantiated in [386].
Finally, the evolution of spread complexity following local quench protocols in 2D CFTs was studied in [155]. The initial state is given by two thermal states on semi-infinite lines joined together at $x=0$ and evolved with the CFT Hamiltonian. The crucial object, as discussed in Sec. 6, is the return amplitude that reads
$S(t)=\left(\frac{\sinh\left(\frac{\pi(t+2{\rm i}\varepsilon)}{\beta}\right)}{\sinh\left(\frac{2\pi{\rm i}\varepsilon}{\beta}\right)}\right)^{-c/8}\,,$ (137)
where $\beta$ is the inverse temperature, $\varepsilon$ is the UV regulator and $c$ is the central charge of the 2D CFT. Interestingly, the return amplitude leads to Lanczos coefficients that depend on the central charge
$a_{n}=\frac{2\pi}{\beta\tan\left(\frac{2\pi\varepsilon}{\beta}\right)}\left(n+\frac{c}{16}\right),\qquad b_{n}=\frac{\pi}{\beta\sin\left(\frac{2\pi\varepsilon}{\beta}\right)}\sqrt{n\left(n+\frac{c}{8}-1\right)}\,.$ (138)
These are examples of the $\mathbb{R}$) analytical solutions discussed in Sec. 6.3.1. We can see that, similarly to the SYK example, the large–$c$ limit effectively constrains the range of $n$. Indeed, if we first extract the naive large–$c$ limit, Lanczos coefficients resemble the Heisenberg-Weyl example with $b_{n}\sim\sqrt{n}$ and quadratic growth of complexity. This is correct as long as $n\ll c/8$. If $n$ is comparable or larger than $c/8$, we have to take the full answer first.
This can be seen from the full answer for the spread complexity which is proportional to the central charge of the CFT as follows
@@
It grow quadratically only for early times, but at late times it grows exponentially.
This local quench setup is described holographically using the AdS/BCFT correspondence [387], where the end-of-the-world (EOW) brane is time-dependent. Interestingly, the tip of the EOW brane, which probes the deepest part of the AdS bulk, follows a geodesic of a massive particle (with heavy mass $m=c/16$). The radial momentum of this particle, computed in the proper distance coordinate, precisely satisfies
$\partial_{t}\mathcal{C}_{K}(t)=-\frac{1}{2\varepsilon}P_{\rho}(t).$ (140)
These explicit checks suggest that, depending on the initial state, it may be possible to engineer precise holographic setups where the spread complexity rate of change can be directly matched with a quantity in gravity and then integrated. The overall dependence on the central charge in (139) is particularly encouraging and enters similarly as in the single-interval entanglement entropy in 2D CFTs that can be reproduced from the Ryu-Takayanagi formula [5].
To finish, let us point out that an intriguing perspective on the Krylov basis was also presented in [156], where, using Wigner functions, the authors argued that the Krylov basis is ”the most classical” basis in Quantum Field Theory, potentially making it a natural candidate for describing the evolution of semi-classical states in the bulk code subspace.
7.6 Nielsen’s complexity and holography
There are also attempts to understand holographic complexity by exploiting equivalence between bulk and boundary symplectic forms [388, 48, 30], inspired by behavior under conformal transformations in Bañados geometries [136, 389]. [^38] Alternative attempts, where gravity arises as a consequence of spacetime optimizing the computational cost of its own dynamics, are discussed in [390, 41, 391].
The relation to symplectic forms also motivated the connection between quantum complexity in CFTs and gravity for the Fubini-Study cost function obtained in section 5.6.3. To understand this relation, we observe that one can map trajectories built from the unitary operators (56) of a CFTd to the geodesics of a massive particle of mass $m$ in AdSd+1 spacetime. The Fubini-Study metric (57b) receives a natural geometric interpretation in terms of the minimal $\delta X_{\rm min}$ and maximal $\delta X_{\rm max}$ perpendicular distances between infinitesimally nearby geodesics (as illustrated in Fig. 1 of Ref. [42])
@@ 7.6 Nielsen’s complexity and holography
It would be interesting to generalize the relation (141) to complexity associated with states that are related by a finite transformation. Furthermore, the connection to cost functions other than the Fubini-Study one is a subject of an ongoing investigation [140].
@@
For general excited states in holography, we must deal with two distinct bulk geometries whose direct comparison should occur via covariant notions (for example, through one of the holographic complexity proposals) associated with each of the corresponding asymptotic boundaries. The authors of  [44] proposed the following idea. For circuits in which the circuit parameter is the physical time in QFT, one layer of the circuit is simply given by the instantaneous QFT Hamiltonian. This Hamiltonian is defined in terms of the energy-momentum tensor components smeared along a time slice, e.g. $H=\int d^{d-1}x\,T_{tt}(\vec{x})$. The infinitesimal increment of the Fubini-Study distance (57b) is proportional to the variance of the Hamiltonian in an instantaneous state. However, given that the Hamiltonian is a sum of local operators, its variance is given by a sum of two-point functions of local operators. In AdS/CFT, the holographic dictionary [365, 366, 392, 393] provides a systematic geometric way of computing correlation functions of the boundary QFTs. As a result, the Fubini-Study cost (57b) associated with one layer of time evolution on the boundary is always geometric by the virtue that all its ingredients are geometric. This statement therefore expresses an exact mapping between two geometries: the auxiliary complexity geometry of the Fubini-Study cost on one hand, and the gravitational geometry of the bulk on the other. While the above statement is geometric, the bulk manifestation of the Fubini-Study cost that leads to it is rather complicated and implicit. However, in the context of holographic two-dimensional CFTs and states represented holographically as solutions of the Einstein’s equations with negative cosmological constant in three bulk dimensions, one can relate two-point functions of the energy-momentum tensor to geodesic lengths in the bulk [44]. This leads to a much more explicit expression for the Fubini-Study cost and a direct connection to the kinematic space program [394, 395, 396].